# Application-Driven Value Alignment in Agentic AI Systems: Survey and Perspectives

### v1

@@ Abstract
The ongoing evolution of AI paradigms has propelled AI research into the Agentic AI stage. Consequently, the focus of research has shifted from single agents and simple applications towards multi-agent autonomous decision-making and task collaboration in complex environments. As Large Language Models (LLMs) advance, their applications become more diverse and complex, leading to increasingly situational and systemic risks. This has brought significant attention to value alignment for AI agents, which aims to ensure that an agent’s goals, preferences, and behaviors align with human values and societal norms. This paper reviews value alignment in agent systems within specific application scenarios. It integrates the advancements in AI driven by large models with the demands of social governance. Our review covers value principles, agent system application scenarios, and agent value alignment evaluation. Specifically, value principles are organized hierarchically from a top-down perspective, encompassing macro, meso, and micro levels. Agent system application scenarios are categorized and reviewed from a general-to-specific viewpoint. Agent value alignment evaluation systematically examines datasets for value alignment assessment and relevant value alignment methods. Additionally, we delve into value coordination among multiple agents within agent systems. Finally, we propose several potential research directions in this field.
@@
Value Alignment, LLM-based Agent System Application, Agentic AI Systems, AI Agent.
@@ 1 Introduction
[]The rapid advancement of Large Language Models (LLMs) in recent years has profoundly transformed human society’s production and lifestyle. Leveraging deep learning techniques and vast training datasets, LLMs such as OpenAI’s GPT series and Google’s Gemini series have demonstrated exceptional natural language understanding and generation capabilities [1, 2]. Utilizing LLMs as their ’brains’, LLM-based agents have found extensive applications across various knowledge-intensive domains, including content creation, automated programming, education, healthcare, and business decision-making. Their increasing ability to autonomously participate in decision making, collaboration, and governance has significantly improved efficiency and provided unprecedented interactive experiences [3]. Furthermore, researchers are increasingly exploring the coordination of multi-agent systems to achieve collective intelligence. AI agents are increasingly evolving towards autonomous systems, forming a new paradigm: Agentic AI. Characterized by multi-agent collaboration, dynamic task decomposition, persistent memory, and autonomous decision-making capabilities, this paradigm is becoming a central focus for both academia and industry [4].
The evolution of this new AI paradigm urgently requires researchers to move beyond the limitations of single-agent perspectives and isolated application domains. We need to explore the deep coupling between agent systems and their application scenarios. As LLMs continue to break through, the tasks they perform and the scenarios in which they are applied become more diverse and complex, and the societal risks surrounding them gradually increase. LLMs may retain harmful information in training data, disclose private data, generate misleading information, and even perform behaviors that are harmful to humans [5]. The various undesirable behaviors, such as manipulation and deception, exhibited by LLM-based agent systems [6] have raised concerns about the possible ethical and safety challenges posed by AI systems. If future AI development surpasses human cognitive capabilities and acquires autonomous learning and decision-making abilities, the goal systems it establishes may deviate from or even contradict the original intentions of human designers. In autonomously pursuing their objectives, such agents could mobilize substantial resources to sustain their operations and continuously improve their capabilities. This would systematically displace humans in terms of resource allocation and decision-making authority. Without effective control mechanisms and robust value alignment safeguards, this dynamic is likely to lead to a loss of human dominance in decision-making systems, ultimately triggering a catastrophic existential crisis [7]. As LLMs become increasingly embedded in specific application scenarios, their associated risks are gradually acquiring situational and systemic characteristics. For example, in multi-agent systems used for urban traffic optimization, the agents may prioritize efficiency over fairness, thereby exacerbating social inequality [8]. In the healthcare sector, LLMs may provide diagnostic or therapeutic recommendations without sufficient clinical evidence or by misinterpreting ethical guidelines, potentially resulting in serious consequences such as breaches of patient privacy, misdiagnoses, or overtreatment [9]. Consequently, ensuring that LLMs adhere to human instructions and values while avoiding harmful behaviors has become an urgent challenge.
Moreover, as AI systems represented by LLMs become increasingly integrated into governance scenarios, how to effectively manage the institutional frictions and transaction costs they generate has become a critical practical and theoretical issue. The emergence of multi-agent systems has led to the deployment of multiple LLMs, each possessing autonomous language comprehension and generation capabilities, within collaborative environments. In such settings, these LLMs work together to accomplish tasks involving decision-making, planning, and execution. The deployment and application of LLMs in this context typically entail intricate mechanisms for information sharing, access control, and collaboration among diverse agents. While this facilitates flexible cooperation, it simultaneously introduces unprecedented challenges related to value conflicts, heterogeneous objectives, and unpredictable agent behaviors. Without sound institutional arrangements and well-designed interaction mechanisms, complex organizational structures and ambiguous task boundaries can lead to low coordination efficiency, unclear responsibilities, and a lack of trust mechanisms, thereby increasing governance costs and undermining administrative effectiveness [10].
@@
Survey Year Keywords Research Subject Value Principles Application Scenarios Value Alignment Datasets Value Alignment Methods Value Coordination
@@
Ours 2025 Value alignment principles, dataset curation, technical methods, and value coordination for LLM-based agents Single-Agent, Multi-agent systems
 [1] 2023 Communication structures, practical applications and societal systems etc. of LLM-based agents Single-Agent, Multi-agent systems - - - - -
 [11] 2025 Value alignment objectives, datasets, techniques, and evaluation methods for LLM-based agents Multi-agent systems - - - - -
 [12] 2025 Conceptual framework, interaction mechanisms, and application overview of LLM-based agent systems Multi-agent systems - - - - -
 [13] 2024 Capabilities, framework Analysis, and application over- view of LLM-based multi-agent systems Multi-agent systems - - - - -
 [14] 2024 Constituent modules, application overview, and eval- uation methods of LLM-based autonomous agents Agents - - - - - - -
 [15] 2023 Motivations and objectives, alignment methods, and assurance and governance of AI alignment LLMs, Agents - - - -
 [11] 2023 Definition and evaluation of LLM alignment objectives LLMs - - - - -
 [16] 2023 Definition, categories, testing, and evaluation of LLM alignment LLMs - - - - -
 [17] 2023 Definition, normative principles, and technical methods of LLM value alignment LLMs - - - - -
 [18] 2023 Alignment objectives for trustworthy LLMs LLMs - - - - - -
 [19] 2024 Challenges, fundamental definitions, and alignment frameworks of LLM value alignment LLMs - - - - -
 [20] 2025 Necessity, conceptual definitions, theoretical approaches, challenges and future outlook of LLM value alignment LLMs - - - - - -
@@
However, these detrimental behaviors and institutional frictions fundamentally arise from deep-seated tensions among multiple stakeholders concerning governance objectives, value concepts, and behavioral norms. Effectively embedding LLMs within human value systems is a critical prerequisite for responsibly unleashing their potential. Value alignment for AI agents—ensuring that their goals, preferences, and behavioral outputs during operation are consistent with the core values of individuals, groups, or institutions—is increasingly recognized as a solution to these issues [21]. In single LLM agent scenarios, alignment techniques primarily focus on guiding models to generate content that adheres to ethical norms and societal expectations through methods such as human feedback and supervised fine-tuning, thereby avoiding harmful outputs, misinformation, and moral deviations [8]. In complex multi-agent systems, value alignment becomes an organizational challenge concerning how multiple agents develop shared norms and coordination mechanisms. Therefore, from a governance perspective, solutions should not be limited to value alignment within single agents but also explore complex value alignment within multi-agent systems’ interaction mechanisms and organizational structures. Value alignment in multi-agent systems is no longer merely an isolated algorithmic optimization problem; it has become a foundational issue concerning agent behavioral boundaries and system stability. Since Agentic AI is fundamentally characterized by scenario-driven multi-agent coordination, its core research must focus on task decomposition mechanisms, interaction design, and environmental adaptability strategies for multiple agents within specific application contexts. Furthermore, understanding the interaction mechanisms and organizational patterns of multi-agent systems in complex environments is crucial for exploring Agentic AI’s value alignment and ensuring its robust development.
@@
Existing research has surveyed key methods and technological advancements in LLM value alignment  [1, 11, 12, 13, 14, 15, 11, 16, 17, 18, 19, 20], see Table I. However, few studies have specifically addressed value alignment within LLM-based multi-agent systems. While some studies have explored multi-agent coordination mechanisms, a comprehensive investigation into value alignment remains notably absent. This gap is particularly pronounced in LLM-based multi-agent systems, where their inherent high autonomy and language-driven collaborative modes further amplify issues of value heterogeneity, goal inconsistency, and fragile cooperation mechanisms. Furthermore, most existing research largely remains at theoretical or general algorithmic levels. There’s limited focus on the mechanisms and strategies for multi-agent value alignment in specific application scenarios, alongside a lack of hierarchical categorization for value alignment within these systems. This suggests that value alignment shouldn’t be understood merely as a technical task at the model level. Instead, it must be viewed as a systemic governance issue that permeates the design of interaction mechanisms and the configuration of organizational structures.
@@
This paper reviews value alignment in LLM-based agent systems for various application scenarios, integrating insights from the evolution of AI technology and the demands of social governance, and proposes future outlooks. As a research review, Sections 2, 3, and 4 comprehensively survey value principles, agent system application scenarios, and agent value alignment evaluation. Specifically, value principles are organized hierarchically from a top-down perspective, encompassing macro, meso, and micro levels. Intelligent agent application scenarios are categorized and reviewed from a general-to-specific viewpoint. Agent value alignment evaluation systematically examines datasets for value alignment assessment and the methods used for agent value alignment. Section 5 presents insights and future outlook. We advocate for the establishment of a multi-level value alignment evaluation system. We also explore value coordination among multiple agents within intelligent agent systems, focusing particularly on values in the design of inter-agent interaction mechanisms and organizational models. Furthermore, we propose that to better facilitate value alignment for agents in specific scenarios, enterprise organizations could share value propositions from datasets in micro-level domains, provided no confidential information is involved. This paper bridges the gap between theoretical alignment research and the practical demands of deploying LLM agent systems in real-world applications. It offers new research perspectives on LLM agent system value alignment by considering actual application scenarios and organizational interaction dynamics, thereby advancing the exploration of Agentic AI’s internal mechanisms.
2 Agent Value Alignment
@@
This section introduces values and value alignment, and then systematically categorizes specific value principles from a top-down, hierarchical perspective encompassing macro, meso, and micro levels. Before delving into the value alignment challenges of LLM-based agent systems in subsequent sections, we first clarify the core concepts used in this paper, particularly the definition of "values," to prevent any ambiguity in understanding.
@@ 2.1 Value
In the design and governance of AI systems, particularly agents, values serve not only as normative bases guiding behavior generation but also as benchmarks for assessing a system’s social compatibility. These values influence individual behavior, social norms, cultural identity, and collective decision-making processes. However, value itself isn’t a clearly defined or consistently standardized concept; its multifaceted nature and historical evolution contribute to significant heterogeneity.
Values, as inherent guides for human behavior, manifest as a set of stable cognitive principles. These principles not only shape individual judgments about life but also constitute the foundational normative framework for social interaction. They are the fundamental tenets influencing attitudes, choices, and social interactions. Scholars such as [25, 26, 27] have proposed various definitions and theoretical perspectives on values. Kluckhohn was among the first to define value orientations in a theory of action, emphasizing their role in guiding behavior. Rokeach, conversely, pioneered a description linking values with affect, positing that these beliefs guide behavior and the evaluation of people, objects, and events. Schwartz, through systematic cross-cultural comparisons, revealed the universal characteristics and inherent structure of value systems.
Contemporary value research primarily focuses on three core dimensions. First, the systemic characteristics of values are examined; for instance, Schwartz’s theory classifies values into dimensions such as openness to change versus conservation, and self-enhancement versus self-transcendence [28]. Second, measurement methodologies are a key area, with standardized scales widely adopted for quantitative assessment [29]. Finally, the functional research domain explores values’ predictive efficacy for social activities, including group decision-making and environmental behavior. Research on values also investigates their influence on social attitudes, prejudice, political behavior, environmental concerns, and intergroup relations. For example, social value orientations affect cooperative and competitive behavior in negotiation and group contexts. Studies also examine values’ functions in self-image maintenance, moral reasoning, and cultural variations [30].
The role of values in society is fundamental. They regulate social order by shaping internalized norms and enable cultural continuity through intergenerational transmission. Values influence education, social compassion, political attitudes, and pro-environmental intentions [31]. Values are essential for comprehending human motivation, moral judgment, and interpersonal dynamics. Essentially, values serve as cognitive frameworks that guide how individuals and groups set priorities, manage conflicts, and build meaningful social systems [32].
One of the most challenging characteristics of human values lies in their inherent contradictions and dynamic nature. Tensions often exist between values. Therefore, a value system is better understood as a context-dependent, flexible structure rather than a rigid hierarchy [31, 33]. Synthesizing existing research, we contend that human values are not fixed entities but dynamically constructed concepts. When values serve as an alignment objective, it implies LLMs must apply these principles to guide their behavior, thereby maximizing their potential. In this context, the categorization and definition of values become particularly crucial.
This paper broadens the definition of values, considering them as dynamic concepts characterized by a multi-layered structure and generalized agency. First, as both human values and LLM-based agents are still evolving, the multi-agent systems they form exhibit diverse configurations across various application scenarios. We propose that values will be continuously interpreted and reshaped through practice and interaction. Second, aligning with different application scenario tiers, we conceptualize values as a multi-layered system: universal human values, national and industry-level values, and organizational business-level values. This paper does not delve into individual-level value issues. Universal human values are those broadly accepted by humanity. National and industry-level values are reflected in national strategies, policy directives, and industry regulations. Organizational business-level values, conversely, relate to an organization’s distinct culture and operational characteristics. These value layers are nested, with higher-level values constraining and guiding those below. Furthermore, this paper emphasizes that humans aren’t the sole bearers of values. This means that beyond simply injecting values into agents through human programming, agents can also internalize value concepts during their design and interaction processes. Values can be concretized through mechanisms like agent interaction and subsequently transmitted and evolve within the resulting organizational patterns. For instance, a decentralized architecture formed by task- and role-based multi-agent systems might inherently embody values opposing power monopolies and promoting anti-fragility, along with their associated governance orientations. In essence, values are reflected not only in "what agents are endowed with," but also in "how they operate through interaction rules" and "what organizational models they inhabit."

### v2

# Multi-level Value Alignment in Agentic AI Systems: Survey and Perspectives

@@ Abstract
The ongoing evolution of AI paradigms has propelled AI research into the agentic AI stage. Consequently, the focus of research has shifted from single agents and simple applications towards multi-agent autonomous decision-making and task collaboration in complex environments. As Large Language Models (LLMs) advance, their applications become more diverse and complex, leading to increasing situational and systemic risks. This has brought significant attention to value alignment for agentic AI systems, which aims to ensure that an agent’s goals, preferences, and behaviors align with human values and societal norms. Addressing socio-governance demands through a Multi-level Value framework, this study comprehensively reviews value alignment in LLM-based multi-agent systems as the representative archetype of agentic AI systems. Our survey systematically examines three interconnected dimensions: First, value principles are structured via a top-down hierarchy across macro, meso, and micro levels. Second, application scenarios are categorized along a general-to-specific continuum explicitly mirroring these value tiers. Third, value alignment methods and evaluation are mapped to this tiered framework through systematic examination of benchmarking datasets and relevant methodologies. Additionally, we delve into value coordination among multiple agents within agentic AI systems. Finally, we propose several potential research directions in this field.
@@
Multi-level Value Alignment, Agentic AI Systems, AI Agent, LLM-based Agent System Application.
@@ 1 Introduction
The rapid advancement of Large Language Models (LLMs) in recent years has profoundly transformed human society’s production and lifestyle. Leveraging deep learning techniques and vast training datasets, LLMs have demonstrated exceptional natural language understanding and generation capabilities [1, 2]. Utilizing LLMs as their ’brains’, LLM-based agents have found widespread applications in knowledge-intensive domains, including content creation, automated programming, education, healthcare, and business decision-making. Their increasing ability to autonomously participate in decision making, collaboration, and governance has significantly improved efficiency and provided unprecedented interactive experiences [3]. Furthermore, researchers are increasingly exploring the coordination of multi-agent systems to achieve collective intelligence. AI agents are increasingly evolving towards autonomous systems, forming a new paradigm: Agentic AI. Characterized by multi-agent collaboration, dynamic task decomposition, persistent memory, and autonomous decision-making capabilities, agentic AI is becoming a central focus for both academia and industry [4]. Without loss of generality, we use the terms “AI Agent”, “agent systems”, “LLM-based agent systems”, “multi-agent systems” interchangeably to refer to “Agentic AI systems”.
@@
Survey
Year
Keywords
Research Subject
Value Principles
Application
Scenarios
Value Alignment
Datasets
Value Alignment
Methods
Value Coordination
@@ Methods
Ours 2025
Value alignment principles, dataset curation, technical
methods, and value coordination for LLM-based agents
Single-Agent,
Multi-agent systems
 [1] 2025
Communication structures, practical applications and
societal systems etc. of LLM-based agents
Single-Agent,
Multi-agent systems
 [5] 2025
Value alignment objectives, datasets, techniques, and
evaluation methods for LLM-based agents
Multi-agent systems - - - - -
 [6] 2025
Conceptual framework, interaction mechanisms, and
application overview of LLM-based agent systems
Multi-agent systems - - - - -
 [7] 2024
Capabilities, framework Analysis, and application over-
view of LLM-based multi-agent systems
Multi-agent systems - - - - -
 [8] 2024
Constituent modules, application overview, and eval-
uation methods of LLM-based autonomous agents
Agents - - - - - - -
 [9] 2023
Motivations and objectives, alignment methods, and
assurance and governance of AI alignment
LLMs, Agents - - - -
 [10] 2023
Definition and evaluation of LLM alignment objectives
LLMs - - - - -
 [11] 2023
Definition, categories, testing, and evaluation of LLM alignment
LLMs - - - - -
 [12] 2023
Definition, normative principles, and technical methods
of LLM value alignment
LLMs - - - - -
 [13] 2023
Alignment objectives for trustworthy LLMs
LLMs - - - - - -
 [14] 2024
Challenges, fundamental definitions, and alignment
frameworks of LLM value alignment
LLMs - - - - -
 [15] 2025
Necessity, conceptual definitions, theoretical approaches,
challenges and future outlook of LLM value alignment
LLMs - - - - - -
The evolution of this new AI paradigm urgently requires researchers to move beyond the limitations of single-agent perspectives and isolated application domains. We need to explore the deep coupling between agent systems and their application scenarios. As LLMs continue to break through, the tasks they perform and the scenarios in which they are applied become more diverse and complex, and the societal risks surrounding them gradually increase. LLMs may retain harmful information in training data, disclose private data, generate misleading information, and even perform behaviors that are harmful to humans [16]. The various undesirable behaviors, such as manipulation and deception, exhibited by LLM-based agent systems [17] have raised concerns about the possible ethical and safety challenges posed by AI systems. As LLMs become increasingly embedded in specific application scenarios, their associated risks are gradually acquiring situational and systemic characteristics. For example, in multi-agent systems used for urban traffic optimization, the agents may prioritize efficiency over fairness, thereby exacerbating social inequality [18]. In the healthcare sector, LLMs may provide diagnostic or therapeutic recommendations without sufficient clinical evidence or by misinterpreting ethical guidelines, potentially resulting in serious consequences such as breaches of patient privacy, misdiagnoses, or overtreatment [19].
As agentic AI systems increasingly integrate into diverse governance scenarios, controlling the resulting institutional friction and transaction costs during their operation and public governance becomes a critical practical and theoretical concern. In the evolving landscape of agentic AI systems, multiple LLMs with autonomous language understanding and generation capabilities are embedded in collaborative environments to perform decision-making, planning, and execution tasks. Such deployments typically involve information sharing, permission configuration, and coordination mechanisms among diverse stakeholders. While enabling flexible collaboration, this also introduces unprecedented challenges like value conflicts, heterogeneous objectives, and unpredictable behaviors. Without robust institutional arrangements and well-designed interaction mechanisms, complex organizational structures and fluid task boundaries can lead to inefficient coordination, blurred accountability, and a lack of trust, consequently escalating governance costs and undermining executive capacity [20].
@@
However, these detrimental behaviors and institutional frictions fundamentally stem from deep-seated tensions among multiple stakeholders concerning governance objectives, value concepts, and behavioral norms. Effectively embedding human value systems into LLMs is a critical prerequisite for responsibly unleashing LLMs’ potential. Value alignment for AI agents, ensuring that their goals, preferences, and behavioral outputs during operation align with the core values of individuals, groups, or institutions, is increasingly recognized as a solution to these issues [21]. In single LLM agent scenarios, alignment techniques primarily focus on guiding models to generate content that adheres to ethical norms and societal expectations via human feedback and supervised fine-tuning, thereby avoiding harmful outputs, misinformation, and moral deviations [18]. In complex multi-agent systems, value alignment becomes an organizational challenge concerning how multiple agents develop shared norms and coordination mechanisms. Therefore, from a governance perspective, solutions should not be limited to value alignment within a single agent but also explore complex value alignment within multi-agent systems’ interaction mechanisms and organizational structures. Value alignment in agentic AI systems is no longer merely an isolated algorithmic optimization problem. Instead, it is a foundational issue that dictates agent behavioral boundaries and system stability. It represents a complex undertaking involving a delicate balance among diverse interests, societal values, and practical utility. Since agentic AI is fundamentally characterized by scenario-driven multi-agent coordination, its core research must focus on task decomposition mechanisms, interaction design, and environmental adaptability strategies for multiple agents within specific application contexts. Furthermore, focusing on the interaction mechanisms and organizational patterns of multi-agent systems in complex environments is crucial for exploring agentic AI’s value alignment and robust development, ultimately serving human well-being.
@@
Existing research has surveyed methods and technological advancements in LLM value alignment, see Table I. However, few studies have specifically addressed value alignment within LLM-based multi-agent systems. While some studies have explored multi-agent coordination mechanisms, a comprehensive investigation into value alignment remains notably absent. Furthermore, most existing research largely remains at theoretical or algorithmic levels. There is limited focus on the mechanisms and strategies for multi-agent value alignment in specific application scenarios, alongside a lack of hierarchical categorization for value alignment within these systems. This suggests that value alignment should not be understood merely as a technical task at the model level. Instead, it must be viewed as a systemic governance issue that permeates the design of interaction mechanisms and the configuration of organizational structures.
Building on these gaps, this study comprehensively reviews value alignment in LLM-based multi-agent systems as the representative archetype of agentic AI systems. From a socio-governance perspective, we synthesize over 200 publications to survey value principles, application scenarios, value alignment methods and evaluation approaches within a multi-level value framework. We further propose future research directions for value coordination in agentic AI.
@@
To be specific, Sections 2, 3, and 4, as illustrated in Figure 1, comprehensively survey value principles, agent system application scenarios, and agent value alignment evaluation. Specifically, value principles are organized hierarchically from a top-down perspective, encompassing macro, meso, and micro levels. Intelligent agent application scenarios are categorized and reviewed from a general-to-specific viewpoint. Agent value alignment evaluation systematically examines datasets for value alignment assessment and the methods used for agent value alignment. Section 5 presents future directions. We discuss multi-agent value coordination within agent systems, with a particular focus on the design of agent interaction mechanisms and organizational models. We also present our outlook on value alignment issues in communication protocols and advocate for the establishment of a multi-level value alignment evaluation system and the collaborative construction of open datasets with society.
To the best of our knowledge, our survey is the first comprehensive review examining value alignment in LLM-based multi-agent systems from the perspective of societal governance and multi-level value structures. It bridges the gap between theoretical alignment research and the practical demands of deploying such systems in real-world applications, offering a novel framework for understanding value alignment in complex autonomous systems.
2 Application-oriented Value Alignment of Agentic AI Systems
@@
This section first introduces our review subjects, agentic AI systems, values and value alignment, and then systematically categorizes specific value principles from a top-down, hierarchical perspective encompassing macro, meso, and micro levels. Before delving into the value alignment challenges of agentic AI systems in subsequent sections, we first clarify the core concepts used in this paper, particularly the definition of “values,” to avoid ambiguity.
2.1 Agentic AI Systems
In this study, we focus on agentic AI systems driven by LLMs. LLMs, such as Deepseek, GPT-4, GPT-3, and ChatGPT, are deep learning models trained on large-scale corpora, capable of understanding and generating natural language text. Advances in model architectures, pretraining corpora, and alignment techniques have significantly enhanced LLMs’ capabilities in knowledge acquisition, instruction following, planning, and reasoning [1], establishing a solid foundation for constructing intelligent digital agents [22]. Serving as a central controller, an LLM can endow an agent with decision-making and memory capabilities [23]. This shift marks an evolution in the AI paradigm—from static, task-specific models to dynamic, interactive agent architectures suitable for complex environments.
Before the rise of LLMs, multi-agent systems (MAS) had already emerged. MAS is a computational system of multiple agents operating and interacting in a shared environment [6]. Enabled by natural language and custom content [24], these agents autonomously collaborate to solve complex problems or complete tasks [25]. The core components of MAS include agents, environment, interaction rules, and organizational patterns [6]. MAS interactions resemble human division of labor and aim to achieve collective intelligence. Each agent holds task-specific knowledge and goals. Powered by LLMs’ language understanding and generation, agents engage in adaptive collaboration for planning, discussion, decision-making, and mutual learning [25, 23, 6], ultimately acting on environmental perception and understanding.
MAS exhibit significant flexibility, robustness, self-organization, coordination, and real-time operation capabilities [6]. While interactions are governed by predefined rules, new behaviors and phenomena may still emerge [26]. Task decomposition and resource sharing in MAS reduce the computational burden on individual agents, making them more cost-effective than single-agent systems, particularly in complex, dynamic tasks that require diverse strategies and collective decision-making [23].
@@ 2.2 Value
As intrinsic guides of human behavior, values represent a stable set of cognitive principles that serve as the foundational norms for social interactions. These principles shape individual attitudes, choices, interpersonal behaviors and group dynamics, as well as serve as fundamental constructs for explaining how individuals and groups prioritize goals, resolve conflicts, and build meaningful social systems [27]. In AI systems, particularly in agent design and governance, values serve both as normative principles guiding behavior and as standards for assessing the social alignment of these systems. Values influence agents’ actions, social norms, cultural identity, and collective decision-making. Understanding values is thus critical to uncovering human motivations, moral judgments, and mechanisms of social interaction.
One of the most challenging aspects of human values is their intrinsic contradiction and dynamic nature. Rather than forming a rigid hierarchical structure, value systems are elastic and context-dependent [28, 29]. Synthesizing previous studies, we conceptualize human values as dynamically constructed notions. When values become alignment targets, LLMs must apply value principles to regulate their own behaviors, thereby maximizing their potential.
This paper expands the conceptualization of values, defining them as dynamic value constructs characterized by a multi-level structure and generalized agency. Both human values and LLM-based agents are evolving entities, thus MAS form exhibit diverse configurations across application scenarios. We argue that values are continuously interpreted and reshaped through practice and interaction. Aligned with different layers of application contexts, values are presented as a multi-tiered conceptual system comprising three levels: universally held human values, values at the national and industry levels, and values specific to organizational operations. This paper does not address individual-level values. These value layers are nested, with higher-level values constraining and guiding lower-level ones.
Furthermore, we emphasize that values are not solely human-borne. Beyond being explicitly embedded by human designers, agents can also “internalize” value principles through design and interaction. Values may be instantiated via agent interactions and inherited or transformed within emergent organizational patterns. For example, MAS with role- and task-based structures may yield decentralized architectures, embodying values such as resistance to power monopolies and mitigation of systemic fragility.